OpenStudy (anonymous):

Company A, company B, company C are the top three cell phone providers in a certain area of the country. the ratio of subscribers of A to B to C is 5:4:3. if company A has 1000 subscriptions, what is the total number of subscribers between company A, company B, company C?

OpenStudy (anonymous):

12000

OpenStudy (anonymous):

may i know the method please...:)

OpenStudy (anonymous):

Let the total number of subscribers between company A, company B, company C be x \[\frac{5}{5+4+3} = \frac{1000}{x}\] 5x=12000 x=2400 answer= 2400 customers. IE: No of customers of A/total customers = a ratio =5/(5+4+3)

OpenStudy (anonymous):

thanks alot...:)